Kayann:

Ever since Renata moved to her new home, she's been keeping track of the height of the tree outside her window. H(t)H(t)H, left parenthesis, t, right parenthesis models the height of the tree (in centimeters), ttt years since Renata moved in. H(t)=210+33tH(t)=210+33tH, left parenthesis, t, right parenthesis, equals, 210, plus, 33, t How fast does the tree grow? centimeters per year..........no diagrams

Gdeinward:

Please post a screenshot of the question or something. This is just too much gibberish.

Mercury:

H(t)=210+33t where H(t) is the height and t is the # of years the slope, therefore, will be the change in height per year, or in other words "how fast the tree grows" simply identify the slope of H(t)=210+33t. major hint: compare to y = mx + b form where m is the slope